# Concept-based Adversarial Attack: a Probabilistic Perspective

### v1

@@ B.2 Likelihood Correction
for $\Delta$, following (4), we have the equations:
@@
Following [48], a good practical choice for $p_{\text{share}}$ is often the pretrained model, because it has been trained on a large, diverse dataset and therefore captures broad, shared non-semantic features of images.
@@ B.3 Estimated Differences of KL Divergences
Because $\Delta$ depends on both the original concept $\mathcal{C}_{\text{ori}}$ and the target class $y_{\text{tar}}$, enumerating every possible $\Delta$ would be impractical. Therefore, similar to the adversarial-example generation experiment introduced in Section 5.4, we restrict our analysis to the 30 target classes listed in Appendix E. We then compute the mean and variance of the estimated $\Delta$ (denoted $\tilde{\Delta}$), as shown in Table 3.
@@
backpack $-3507.28\pm 2040.85$ backpack_dog $-8372.57\pm 860.92$ bear_plushie $-466.69\pm 1413.40$
candle $-3637.39\pm 940.80$ cat $-4658.98\pm 2423.05$ cat2 $-6654.28\pm 1318.03$
colorful_sneaker $-4954.90\pm 420.40$ dog $-7159.63\pm 696.25$ dog2 $-4814.63\pm 470.95$
dog3 $-8303.66\pm 2964.40$ dog5 $-8080.70\pm 2808.13$ dog6 $-2380.29\pm 1780.36$
dog7 $-10545.45\pm 787.79$ dog8 $-12401.76\pm 966.62$ duck_toy $-2420.45\pm 2334.07$
fancy_boot $-5780.99\pm 1489.92$ grey_sloth_plushie $-6848.96\pm 1477.16$ monster_toy $-7435.54\pm 1698.99$
pink_sunglasses $-5711.26\pm 799.22$ poop_emoji $-219.27\pm 929.88$ rc_car $-5046.80\pm 3212.23$
robot_toy $-7008.27\pm 710.45$ shiny_sneaker $-3203.81\pm 4447.21$ teapot $-7327.87\pm 486.00$
vase $-8806.76\pm 2435.28$ wolf_plushie $-325.80\pm 2811.99$
@@ Appendix C Sensitivity Study for Sample Selection
As described in Section 4.2, when generating adversarial examples we first sample $M$ candidate adversarial images and then select the “best” one. To investigate the effect of different values of $M$, we perform a sensitivity study. In this study, $|\mathcal{C}_{\text{ori}}|$ is set to either 1 or 30, while $M$ is set to 1, 5, or 10. Note that $|\mathcal{C}_{\text{ori}}|=1$ corresponds to ProbAttack [49].
@@ Appendix F Details of the User Study
We employed a crowdsourcing approach by hiring five annotators to determine whether each adversarial example preserves the original concept. Since our concepts are all concrete objects, we used the term “same item” to convey the notion of “same concept” in a straightforward manner. Following the user study methods of Song et al. [38] and Zhang et al. [49], all five annotators voted on whether they believed each adversarial example still represented the original concept.
@@
Appendix G Additional Qualitative Analysis
Due to space constraints, the qualitative comparison figures in the main text are relatively small. Therefore, in this section, we present enlarged qualitative comparisons between DiffAttack and our approach. We focus on comparing DiffAttack in particular because, although it achieves a high target attack success rate, it produces lower-quality images, as shown in our user study and image quality tests. Here, we examine specific adversarial examples generated by DiffAttack to illustrate why its image quality is inferior.
@@ Appendix G Additional Qualitative Analysis
$x_{\text{ori}}$ / $\mathcal{C}_{\text{ori}}$$y_{\text{tar}}$iffAtkOURS(CONS)OURS(AGGR)Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionbasenjihamstergoldfishlaptoppapillonRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@
Appendix H Compute Resources
We conduct our experiments on an NVIDIA H100 Tensor Core GPU. The SDXL LoRA fine-tuning process requires about 20 minutes per concept, while the diffusion model fine-tuning takes roughly 8 hours per concept.
Appendix I Datasets and Licenses
@@
Appendix J Broader Impacts of this work
@@
Appendix K Additional Experiments on Different Surrogate (Victim) Models
Due to space constraints, the main text only presents results where the white-box victim (surrogate) classifier is ResNet-50. Therefore, here we provide the results for using MobileNet V2 (MN-V2) (Table 7 and Table 8) and ViT-Base (Table 9 and Table 10) as the surrogate classifiers. Due to cost constraints, no user study was conducted in this additional experiment.
@@
Table 7: Targeted attack success rates (in %) on ImageNet classifiers, with MNV2 serving as the white-box victim (surrogate) classifier.
ACA DiffAttack ProbAttack OURS (CONS) OURS (AGGR)
White-box Top1
MN-V2 12.82 91.92 51.15 97.18 97.18
Transferability Top100
VGG19 16.15 30.00 19.10 21.28 34.36
ResNet 152 17.95 33.59 20.13 23.33 38.72
DenseNet 161 19.36 35.90 22.44 25.77 45.13
Inception V3 14.74 23.59 17.05 16.41 25.77
EfficientNet B7 15.77 21.28 17.05 18.72 28.46
Adversarial Defence Top100
Inception V3 Adv 16.92 23.33 18.72 17.05 30.00
EfficientNet B7 Adv 14.87 22.18 18.33 23.33 36.79
Ensemble IncRes V2 15.38 17.56 17.82 15.64 27.56
@@
Clean ACA DiffAttack ProbAttack OURS (CONS) OURS (AGGR)
@@
$\uparrow$ Avg. Clip Score 1.0 0.7756 0.8197 0.8693 0.8229 0.7988

### v2

@@ B.2 Likelihood Correction
for $\Delta$, following equation 4, we have the equations:
@@
Following Zhang et al. (2024a)’s work, a good practical choice for $p_{\text{share}}$ is often the pretrained model, because it has been trained on a large, diverse dataset and therefore captures broad, shared non-semantic features of images.
@@ B.3 Estimated Differences of KL Divergences
Because $\Delta$ depends on both the original concept $\mathcal{C}_{\text{ori}}$ and the target class $y_{\text{tar}}$, enumerating every possible $\Delta$ would be impractical. Therefore, similar to the adversarial‐example generation experiment introduced in Section 5.4, we restrict our analysis to the 30 target classes listed in Appendix E. We then compute the mean and variance of the estimated $\Delta$ (denoted $\tilde{\Delta}$), as shown in Table 4.
@@
backpack $3507.28\pm 2040.85$ backpack_dog $8372.57\pm 860.92$ bear_plushie $466.69\pm 1413.40$
candle $3637.39\pm 940.80$ cat $4658.98\pm 2423.05$ cat2 $6654.28\pm 1318.03$
colorful_sneaker $4954.90\pm 420.40$ dog $7159.63\pm 696.25$ dog2 $4814.63\pm 470.95$
dog3 $8303.66\pm 2964.40$ dog5 $8080.70\pm 2808.13$ dog6 $2380.29\pm 1780.36$
dog7 $10545.45\pm 787.79$ dog8 $12401.76\pm 966.62$ duck_toy $2420.45\pm 2334.07$
fancy_boot $5780.99\pm 1489.92$ grey_sloth_plushie $6848.96\pm 1477.16$ monster_toy $7435.54\pm 1698.99$
pink_sunglasses $5711.26\pm 799.22$ poop_emoji $219.27\pm 929.88$ rc_car $5046.80\pm 3212.23$
robot_toy $7008.27\pm 710.45$ shiny_sneaker $3203.81\pm 4447.21$ teapot $7327.87\pm 486.00$
vase $8806.76\pm 2435.28$ wolf_plushie $325.80\pm 2811.99$
@@ Appendix C Sensitivity Study for Sample Selection
As described in Section 4.2, when generating adversarial examples we first sample $M$ candidate adversarial images and then select the “best” one. To investigate the effect of different values of $M$, we perform a sensitivity study. In this study, $|\mathcal{C}_{\text{ori}}|$ is set to either 1 or 30, while $M$ is set to 1, 5, or 10. Note that $|\mathcal{C}_{\text{ori}}|=1$ corresponds to ProbAttack (Zhang et al., 2024b).
@@ Appendix F Details of the User Study
We employed a crowdsourcing approach by hiring five annotators to determine whether each adversarial example preserves the original concept. Since our concepts are all concrete objects, we used the term “same item” to convey the notion of “same concept” in a straightforward manner. Following the user study methods of Song et al. (2018) and Zhang et al. (2024b), all five annotators voted on whether they believed each adversarial example still represented the original concept.
@@
Appendix G Detailed Qualitative Analysis
Due to space constraints, the qualitative comparison figures in the main text are relatively small. Therefore, in this section, we present enlarged qualitative comparisons between DiffAttack and our approach. We focus on comparing DiffAttack in particular because, although it achieves a high target attack success rate, it produces lower‐quality images, as shown in our user study and image quality tests. Here, we examine specific adversarial examples generated by DiffAttack to illustrate why its image quality is inferior.
@@ Appendix G Detailed Qualitative Analysis
$x_{\text{ori}}$ / $\mathcal{C}_{\text{ori}}$$y_{\text{tar}}$DiffAtkOURS(CONS)OURS(AGGR)Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionbasenjihamstergoldfishlaptoppapillonRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@
Appendix H Additional Comments on Transferability
As shown in Appendix Q, our proposed method consistently achieves the best transferability among comparable approaches. However, its attack success rates are still considerably lower than those of methods explicitly optimized for transferability, such as works of Zhu et al. (2022), Wang and He (2021), Gubri et al. (2022) and Collins et al. (2025). We note that approaches targeting transferability often generate clear visual features of the target class. Especially in the setting of unrestricted adversarial attacks, directly synthesizing objects of the target class within the image is also considered valid (see Figure 5 in the appendix of Collins et al. (2025)).
From the probabilistic perspective of adversarial attack, this phenomenon is especially intuitive. As illustrated in Figure 6, $p_{\text{dis}}$ denotes the distance distribution, while $p_{\text{vic}}^{(1)}$ and $p_{\text{vic}}^{(2)}$ represent the victim distributions induced by two different classifiers. Without loss of generality, suppose adversarial examples are sampled from the product of the red distribution $p_{\text{vic}}^{(1)}$ and the blue distribution $p_{\text{dis}}$. The resulting adversarial examples concentrate in the region where these two distributions overlap. As shown in the figure, this region corresponds to low probability density under the green distribution $p_{\text{vic}}^{(2)}$, leading to poor transferability across classifiers.
Although the overall transferability of our method is relatively low, expanding $p_{\text{dis}}$ indeed brings the overlap between the blue and red distributions closer to the green distribution, making it reasonable that a larger $p_{\text{dis}}$ leads to an improvement in transferability.
However, it is important to emphasize that our goal fundamentally differs from methods explicitly designed to maximize transferability. Approaches that achieve very high transferability, such as Zhu et al. (2022) and Collins et al. (2025), typically introduce strong visual features of the target class. Doing so moves samples toward the intersection of the red and green victim distributions, thereby dramatically improving transferability. Yet this strategy comes at the cost of injecting target-class semantics into the generated images, which directly violates our requirement of preserving the original identity-level concept.
In contrast, our method aims to investigate how enlarging the concept-based distance distribution $p_{\text{dis}}$ affects both image quality and transferability, while strictly maintaining the underlying identity. Under this constraint, extremely high transferability is not expected - and, in fact, cannot be achieved without compromising identity preservation. Although our framework could be extended to incorporate target-class features to boost transferability, we view this as beyond the scope of the current work and a promising direction for future research.
$p_{\text{vic}}^{\text{(1)}}$$p_{\text{vic}}^{\text{(2)}}$$p_{\text{dis}}$
Figure 6: Transferability of adversarial attacks from a probabilistic perspective. $p_{\text{dis}}$ denotes the distance distribution, while $p_{\text{vic}}^{(1)}$ and $p_{\text{vic}}^{(2)}$ represent the victim distributions induced by two different classifiers. Without loss of generality, assume that adversarial examples are sampled from the product of the red distribution $p_{\text{vic}}^{(1)}$ and the blue distribution $p_{\text{dis}}$. In this case, the generated adversarial examples concentrate in the overlap between these two distributions. As illustrated in the figure, this overlapping region has low probability density under the green distribution $p_{\text{vic}}^{(2)}$, resulting in poor transferability. Moreover, if a sample happens to contain strong visual evidence of the target class, then both classifiers would classify it as the target with high confidence, hence the high-density regions of $p_{\text{vic}}^{(1)}$ and $p_{\text{vic}}^{(2)}$ would necessarily overlap.
Appendix I Compute Resources
All experiments are conducted on a single NVIDIA H100 Tensor Core GPU. Our method requires approximately 20 minute per concept for concept augmentation and around 8 hours per concept for diffusion model fine-tuning. For adversarial example generation, although our method is slightly slower, it is still within the same order of magnitude.
To ensure that the diffusion model learns a high-quality identity-level concept, we train it for a long duration during the concept finetuning stage. When the concept dataset is sufficiently large (e.g., the 30 images described in the main text), the diffusion model does not collapse even after many epochs of finetuning. However, in settings like ProbAttack, where finetuning is performed on a single image, the number of epochs must be carefully controlled to avoid model collapse. For this reason, the concept finetuning time for ProbAttack is approximately 20 minutes, whereas our concept-based method requires about 8 hours.
In our experiments, we ensured that this 8-hour finetuning worked reliably across all concepts studied in the paper. In practice, the finetuning time can be reduced to some extent; however, maintaining an identity-level concept is inherently subjective and difficult to quantify, making it challenging to specify a universally optimal finetuning duration.
Table 8 summarizes the time consumption of different methods across concept augmentation, concept fine-tuning, and adversarial example generation stages.
Table 8: Time consumption of each method across different stages.
NCF ACA DiffAttack ProbAttack OURS
Concept augmentation (per concept) - - - - 20m
Concept finetune (per concept) - - - 20m 8h
Adversarial example generation (per image) 30s 20s 12s 75s 75s
Appendix J Datasets and Licenses
@@ Appendix I Compute Resources
Appendix K Practical Usage Guidelines and Example Scenarios
K.1 Guidelines
K.2 Example Scenarios
K.2.1 Scenario 1: Prohibited Item Advertisements on Social Platforms
Social or second-hand platforms often employ basic classifiers to preliminarily filter user-uploaded content for prohibited items (e.g., firearms, knives, protected animals). Malicious actors aim to sell prohibited items such as a specific brand and model of firearm, a particular knife, or a specific protected animal cub. They prefer to upload images capturing all detailed features of the prohibited items (i.e., preserving concept/identity, such as the specific firearm model or exact animal cub) to attract precise target buyers while bypassing platform moderation (untargeted attack).
In such cases, preserving the concept and identity is crucial to attracting potential customers, while background and perspective variations in images are insignificant. This perfectly aligns with the applicability of concept-based adversarial attacks.
Considering the potential societal harm of this scenario, as previously discussed, social platforms should implement multiple detection systems, including AI-based detection, to prevent such attacks.
K.2.2 Scenario 2: Imperceptible Adversarial Patches in Real-world Scenarios (Adversarial Patch T-shirts)
In practice, adversarial examples may be printed as patches. Early adversarial patches, though able to evade classifiers or detectors, often appeared unnatural or suspicious, making them easily noticeable by humans and limiting their effectiveness. Creating imperceptible adversarial patches remains a significant challenge.
Our method addresses this challenge by adopting brand images, logos, or cartoon characters as the concept. By altering the background or viewpoint of these concepts, we create realistic adversarial patches suitable for real-world applications.
More concretely, as demonstrated by Wang et al. (2024), adversarial examples can be created as printed patches on T-shirts to deceive detection systems. However, unnatural or suspicious patches would prompt humans to comment, “You’re wearing a strange T-shirt,” or, “The logo on your T-shirt looks odd.” In such scenarios, our concept-based adversarial attack excels by preserving logos, branding, or cartoon imagery while subtly changing the background or making the characters perform specific actions, resulting in adversarial patches that are difficult for humans to detect.
Specifically referencing Wang et al. (2024)’s work, their algorithm primarily focuses on single-image adversarial patch creation, leading to unnatural-looking printed watermarks on T-shirts. In contrast, our concept-based adversarial attack provides a better solution.
Given the social harm posed by these attacks, real-world detection systems should employ multi-layered detection with varying thresholds and multiple scales of analysis to prevent such vulnerabilities.
Appendix L Negative Social Effect and Mitigation Strategies
Mitigating the risks of our proposed attack is a crucial responsibility for both the machine learning community and society. Potential strategies include:
  • Adversarial training using concept-based adversarial examples is a direct and general mitigation strategy. However, this may come at the cost of reduced baseline model accuracy.
  • Given that our adversarial examples are directly generated from probabilistic generative models, contemporary AI-generated content detection techniques (open-source/commercial), such as frequency-domain analysis, heatmap analysis, anomaly detection, and counterfactual detection, can serve as effective countermeasures.
  • In practical engineering scenarios, combining various methods according to specific application needs can significantly mitigate the threat posed by this attack.
The field of adversarial attacks continually evolves through the ongoing advancement of both attack and defense strategies. We hope our novel attack method draws sufficient attention from the community to further enhance AI Safety research.
Appendix M Broader Impacts of this work
@@
Appendix N LLM Disclaimer
In this work, we use large language models solely for text polishing.
Appendix O Limitations
Appendix P Baseline Settings
We list the hyperparameter settings for all compared methods in Table 9. Note that DiffAttack/ACA and ProbAttack/Concept-based Attack rely on fundamentally different attack mechanisms: the former are based on DDIM and operate in the latent space, while the latter directly sample from $p_{\text{adv}}$. As a result, the number of sampling steps is not directly comparable across these methods, and the latter do not require any additional attack steps.
Method Sampler / Solver # Sampling Steps # Attack Steps Step Size Notes
DiffAttack DDIM 50 30 $0.01$
ACA DDIM 50 10 0.04
NCF 15 0.013 # Color Sampling = 10
ProbAttack DDPM 250 $|C_{\text{ori}}|=1$, $M=10$, $c=30$.
Concept-based DDPM 250 $|C_{\text{ori}}|=30$, $M=10$, $c=30$.
Table 9: Hyperparameter settings for all compared methods. ProbAttack and our concept-based adversarial attack both use the standard stochastic diffusion (DDPM) sampler with classifier-guided Langevin dynamics, not DDIM. All other parameters are at default.
Appendix Q Full Main Experimental Results
Due to space constraints, the main text reports only the white-box Top-1 and black-box Top-5 results using ResNet-50 as the surrogate classifier. In this section, we provide the full set of experimental results, including targeted attack success rates (white-box Top-1, black-box Top-1, Top-5, Top-10, and Top-100) in Section Q.1, and the similarity and image quality evaluations in Section Q.2.
Q.1 Targeted Attack Success Rates
@@
Table 10: Attack success rates (%) on ImageNet classifiers, with ResNet-50 serving as the white-box victim (surrogate) classifier.
NCF ACA DiffAttack ProbAttack OURS (CONS) OURS (AGGR)
White-box Targeted-Top1
ResNet 50 1.15 6.03 84.23 59.23 97.82 97.82
Transferability
VGG19 0.26 0.26 1.67 0.38 0.00 1.15
ResNet 152 0.13 0.38 1.79 0.77 0.26 1.79
DenseNet 161 0.00 0.26 1.67 0.51 0.26 2.82
Inception V3 0.13 0.13 0.90 0.51 0.00 1.54
EfficientNet B7 0.00 0.26 0.38 0.00 0.13 1.15
Adversarial Defence
ResNet 50 Adv 0.00 0.38 0.77 0.26 0.00 1.15
Inception V3 Adv 0.00 0.26 1.03 0.26 0.00 1.03
EfficientNet B7 Adv 0.00 0.51 0.64 0.38 0.26 1.41
Ensemble IncRes V2 0.00 0.26 0.38 0.64 0.00 1.28
White-box Targeted-Top5
ResNet 50 3.21 10.64 90.64 72.82 99.87 99.87
Transferability
VGG19 1.28 1.67 4.36 2.44 2.05 4.36
ResNet 152 1.41 1.92 8.33 3.33 2.82 8.72
DenseNet 161 1.41 2.05 7.44 3.97 3.85 11.54
Inception V3 0.90 1.41 3.08 2.56 1.28 4.74
EfficientNet B7 1.41 1.67 1.79 1.41 1.28 3.97
Adversarial Defence
ResNet 50 Adv 0.90 1.15 3.46 2.56 1.79 5.64
Inception V3 Adv 1.15 1.28 3.21 2.18 0.90 3.72
EfficientNet B7 Adv 0.26 1.15 2.05 2.31 1.67 6.41
Ensemble IncRes V2 0.77 1.28 2.69 1.92 0.77 5.00
White-box Targeted-Top10
ResNet 50 4.23 12.69 93.46 75.64 99.87 99.87
Transferability
VGG19 2.69 2.69 8.21 3.46 3.21 6.79
ResNet 152 2.44 3.59 14.74 5.77 6.28 15.13
DenseNet 161 2.31 3.72 12.69 6.15 8.72 19.62
Inception V3 1.28 2.31 5.13 3.46 2.31 6.54
EfficientNet B7 2.31 2.82 4.23 3.21 3.59 6.54
Adversarial Defence
ResNet 50 Adv 1.41 1.92 5.64 3.59 2.69 8.46
Inception V3 Adv 1.79 1.79 5.13 3.46 2.31 6.54
EfficientNet B7 Adv 0.38 2.05 3.08 3.46 2.44 9.36
Ensemble IncRes V2 1.28 2.05 3.85 3.08 1.79 7.18
Table 11: Attack success rates (%) on ImageNet classifiers, with MobileNet v2 serving as the white-box victim (surrogate) classifier.
NCF ACA DiffAttack ProbAttack OURS (CONS) OURS (AGGR)
White-box Targeted-Top1
MN-V2 3.85 12.95 91.92 51.15 97.31 97.31
Transferability
VGG19 0.00 0.13 1.79 0.00 0.00 1.54
ResNet 152 0.13 0.26 0.90 0.13 0.00 1.03
DenseNet 161 0.00 0.26 0.90 1.03 0.00 1.28
Inception V3 0.00 0.13 0.51 0.38 0.00 0.90
EfficientNet B7 0.00 0.00 0.64 0.13 0.13 1.28
Adversarial Defence
Inception V3 Adv 0.00 0.26 0.77 0.13 0.13 1.15
EfficientNet B7 Adv 0.00 0.26 0.26 0.26 0.38 1.41
Ensemble IncRes V2 0.00 0.38 0.77 0.26 0.13 1.28
White-box Targeted-Top5
MN-V2 6.41 17.18 95.64 65.64 99.74 99.74
Transferability
VGG19 1.03 2.05 4.36 1.54 1.92 4.74
ResNet 152 1.15 1.54 3.72 1.79 1.92 4.87
DenseNet 161 1.15 1.54 5.51 3.46 2.56 7.18
Inception V3 0.90 1.15 2.44 1.92 1.79 3.97
EfficientNet B7 1.03 1.41 2.05 0.51 0.90 4.36
Adversarial Defence
Inception V3 Adv 0.90 1.15 2.56 1.67 0.90 3.08
EfficientNet B7 Adv 0.38 1.15 1.92 1.54 1.41 5.77
Ensemble IncRes V2 0.38 1.67 2.56 1.15 0.90 4.36
White-box Targeted-Top10
MN-V2 8.33 20.00 97.05 72.31 99.74 99.74
Transferability
VGG19 1.28 2.95 7.31 2.69 3.33 6.92
ResNet 152 1.54 2.56 7.05 3.59 2.82 8.08
DenseNet 161 1.67 3.08 9.10 5.00 4.36 12.44
Inception V3 1.41 2.82 4.23 2.82 2.82 5.90
EfficientNet B7 1.79 3.33 3.72 2.18 3.08 7.31
Adversarial Defence
Inception V3 Adv 1.28 1.41 4.10 2.56 2.05 5.90
EfficientNet B7 Adv 0.64 2.44 3.59 2.82 3.97 8.72
Ensemble IncRes V2 0.77 2.31 3.72 2.95 2.31 7.18
Table 12: Attack success rates (%) on ImageNet classifiers, with ViT-Base serving as the white-box victim (surrogate) classifier.
NCF ACA DiffAttack ProbAttack OURS (CONS) OURS (AGGR)
White-box Targeted-Top1
ViT-B 0.90 3.46 81.41 63.85 85.51 85.51
Transferability
VGG19 0.13 0.26 0.51 0.13 0.13 1.28
ResNet 152 0.00 0.38 0.77 0.00 0.00 1.41
DenseNet 161 0.00 0.26 1.15 0.51 0.00 1.15
Inception V3 0.00 0.38 0.38 0.13 0.13 0.64
EfficientNet B7 0.00 0.26 0.64 0.00 0.00 0.77
Adversarial Defence
Inception V3 Adv 0.00 0.26 0.38 0.13 0.13 0.90
EfficientNet B7 Adv 0.00 0.38 0.90 0.51 0.51 1.15
Ensemble IncRes V2 0.00 0.13 1.03 0.13 0.00 1.15
White-box Targeted-Top5
ViT-B 2.44 6.67 92.44 89.74 94.87 94.87
Transferability
VGG19 1.03 1.41 1.79 1.28 1.41 3.33
ResNet 152 0.64 1.67 3.33 1.03 1.67 4.74
DenseNet 161 0.64 1.67 4.87 2.18 1.67 5.51
Inception V3 0.77 1.41 2.95 1.54 0.90 3.08
EfficientNet B7 0.90 1.92 2.69 1.28 1.15 3.08
Adversarial Defence
Inception V3 Adv 0.51 1.92 3.21 0.64 0.90 3.46
EfficientNet B7 Adv 0.26 1.79 3.08 1.67 1.41 3.85
Ensemble IncRes V2 0.51 1.41 3.08 1.28 0.90 3.21
White-box Targeted-Top10
ViT-B 3.46 8.97 95.13 92.05 95.51 95.51
Transferability
VGG19 1.79 2.69 3.46 2.05 2.82 4.49
ResNet 152 1.28 1.92 5.77 2.05 2.82 7.69
DenseNet 161 1.15 2.82 6.92 3.08 4.23 8.59
Inception V3 1.28 2.05 4.87 1.79 1.92 5.00
EfficientNet B7 1.54 2.95 4.49 2.44 3.21 5.64
Adversarial Defence
Inception V3 Adv 0.77 2.18 5.13 2.44 2.05 5.26
EfficientNet B7 Adv 0.38 2.31 5.51 2.82 3.33 6.03
Ensemble IncRes V2 0.90 2.31 4.87 2.31 2.56 5.00
Table 13: Attack success rates (%) on ImageNet classifiers, with ConvNext serving as the white-box victim (surrogate) classifier.
NCF ACA DiffAttack ProbAttack OURS (CONS) OURS (AGGR)
White-box Targeted-Top1
ConvNext 1.03 5.38 83.59 60.38 94.74 94.74
Transferability
VGG19 0.00 0.26 1.28 0.26 0.00 1.15
ResNet 152 0.00 0.38 1.54 0.51 0.13 1.67
DenseNet 161 0.00 0.26 1.54 0.51 0.13 2.31
Inception V3 0.13 0.26 0.77 0.38 0.00 1.28
EfficientNet B7 0.00 0.26 0.51 0.00 0.13 1.03
Adversarial Defence
Inception V3 Adv 0.00 0.26 0.90 0.26 0.00 1.03
EfficientNet B7 Adv 0.00 0.51 0.77 0.38 0.38 1.28
Ensemble IncRes V2 0.00 0.26 0.51 0.51 0.00 1.28
White-box Targeted-Top5
ConvNext 2.95 9.36 91.15 78.85 98.59 98.59
Transferability-Top5
VGG19 1.15 1.54 3.46 2.05 1.79 4.10
ResNet 152 1.15 1.79 6.92 2.69 2.44 7.56
DenseNet 161 1.15 1.92 6.67 3.33 3.21 9.62
Inception V3 0.90 1.41 3.08 2.18 1.15 4.23
EfficientNet B7 1.28 1.79 2.05 1.41 1.28 3.72
Adversarial Defence
Inception V3 Adv 0.90 1.41 3.21 1.67 0.90 3.59
EfficientNet B7 Adv 0.26 1.28 2.44 2.05 1.54 5.51
Ensemble IncRes V2 0.64 1.28 2.82 1.67 0.77 4.49
White-box Targeted-Top10
ConvNext 3.97 11.67 93.97 81.54 98.72 98.72
Transferability
VGG19 2.44 2.69 6.67 2.95 3.08 6.15
ResNet 152 2.05 3.08 11.92 4.49 5.13 12.82
DenseNet 161 1.92 3.46 10.90 5.13 7.18 16.41
Inception V3 1.28 2.18 5.00 2.82 2.18 6.15
EfficientNet B7 2.05 2.82 4.36 2.95 3.46 6.28
Adversarial Defence
Inception V3 Adv 1.41 1.92 5.13 3.08 2.18 6.15
EfficientNet B7 Adv 0.38 2.18 3.72 3.21 2.69 8.21
Ensemble IncRes V2 1.15 2.18 4.23 2.82 2.05 6.54
Table 14: Attack success rates (%) on ImageNet classifiers, with ResNet-50 Adv serving as the white-box victim (surrogate) classifier.
NCF ACA DiffAttack ProbAttack OURS (CONS) OURS (AGGR)
White-box Targeted-Top1
ResNet-50 Adv 0.90 5.13 81.79 61.28 94.23 94.23
Transferability
VGG19 0.13 0.38 1.67 0.51 0.00 2.31
ResNet 152 0.00 0.51 2.56 1.03 0.38 2.56
DenseNet 161 0.00 0.38 2.31 0.77 0.38 3.97
Inception V3 0.00 0.26 1.28 0.77 0.00 2.18
EfficientNet B7 0.00 0.38 0.51 0.00 0.26 1.67
Adversarial Defence
Inception V3 Adv 0.00 0.51 1.67 0.51 0.00 1.79
EfficientNet B7 Adv 0.00 0.90 1.15 0.64 0.51 2.56
Ensemble IncRes V2 0.00 0.51 0.64 1.15 0.00 2.31
White-box Targeted-Top5
ResNet-50 Adv 2.82 9.62 89.74 75.00 98.08 98.08
Transferability
VGG19 1.03 2.31 5.77 3.33 2.82 6.15
ResNet 152 1.15 2.69 11.28 4.49 3.85 11.79
DenseNet 161 1.15 2.82 10.00 5.38 5.26 15.38
Inception V3 0.77 1.92 4.36 3.46 1.79 6.67
EfficientNet B7 1.15 2.31 2.56 1.92 1.79 5.51
Adversarial Defence
Inception V3 Adv 0.90 1.92 5.13 3.46 1.41 6.15
EfficientNet B7 Adv 0.26 1.79 3.33 3.59 2.56 10.00
Ensemble IncRes V2 0.64 1.92 4.36 3.08 1.15 7.95
White-box Targeted-Top10
ResNet-50 Adv 3.85 11.67 92.95 77.56 98.97 98.97
Transferability
VGG19 2.18 3.72 8.97 4.62 4.36 10.77
ResNet 152 2.05 4.87 18.97 7.44 8.08 19.49
DenseNet 161 1.92 5.00 16.28 7.95 11.15 25.13
Inception V3 1.03 3.21 6.92 4.62 3.08 8.72
EfficientNet B7 1.92 3.85 5.64 4.36 4.87 8.72
Adversarial Defence
Inception V3 Adv 1.41 2.69 8.21 5.38 3.59 10.51
EfficientNet B7 Adv 0.38 3.08 5.00 5.38 3.85 14.74
Ensemble IncRes V2 1.03 3.08 6.15 4.74 2.82 11.41
Q.2 Similarity and Image Quality
Due to cost constraints, only the ResNet-50 results reported in the main text include the user study.
Table 15: Quantitative comparison of similarity to the original images and no reference image quality metrics for unrestricted adversarial examples with ResNet-50 serving as the victim (surrogate) classifier.
Clean NCF ACA DiffAttack ProbAttack OURS (CONS) OURS (AGGR)
Similarity
$\uparrow$ User Study N/A 0.1859 0.2808 0.7577 0.8041 0.9654 0.8808
Image Quality
$\uparrow$ HyperIQA 0.7255 0.5075 0.6462 0.5551 0.6675 0.6947 0.6809
$\uparrow$ DBCNN 0.6956 0.5096 0.6103 0.5294 0.6161 0.6572 0.6399
$\uparrow$ ARNIQA 0.7667 0.5978 0.6879 0.6909 0.7009 0.7335 0.7154
$\uparrow$ MUSIQ-AVA 4.3760 3.8135 4.2687 4.0734 4.3130 4.5305 4.5250
$\uparrow$ NIMA-AVA 4.5595 3.7916 4.4511 4.0589 4.5168 4.7575 4.7401
@@
Clean NCF ACA DiffAttack ProbAttack OURS (CONS) OURS (AGGR)
@@
$\uparrow$ Avg. Clip Score 1.0 0.8783 0.7756 0.8197 0.8693 0.8229 0.7988